OpenStudy (anonymous):

Train A Is going at 84 km/hr. After it went for 100 km, B started to go at 72 km/hr. After 0.5 hours they met. What's the distance in between the 2 trains? They're going towards each other. It has to be in an algebraic formula with X. Please show me the steps, i keep ending up with negatives....

OpenStudy (anonymous):

What grade are you in??

OpenStudy (anonymous):

I'm in 6th.. >.<

OpenStudy (anonymous):

ok! same here i dont get the question sarry

OpenStudy (anonymous):

oh kk, np.

OpenStudy (anonymous):

im really sorry anna

OpenStudy (anonymous):

Well, you know they are at least 100km apart, from the initial motion. Then, after that, if they are moving towards each other at their respective speeds, they are effectively getting closer at the combined speed (72 + 84), and distance = speed x time. If I read the question right, and A doesn't stop after 100km or something.

OpenStudy (anonymous):

Well, this is the formula I got, and the steps. 84+72x=x-100 72x=x-100-84 72x-x=-16 71x=16

OpenStudy (anonymous):

*-16 I'm sure it's not right...

OpenStudy (anonymous):

Where did you get 72x from? I don't really understand what you did, sorry. Total Distance is the amount covered in the whole time. To being with, A moves 100km. Then after, A is moving for 0.5 hours at 84km/h (I think) and B is moving at 72km/h. To total distance is the amount travelled by all of them in this time. 100 + (Distance travelled by A) + (Distance travelled by B) You can work out how far each of them travels by using Distance = speed x time, with 0.5 hours and their respective speeds.